Argonne National Laboratory

Advanced Photon Source
Sector 7

Argonne Home > Advanced Photon Source >

Sector 7 Useful Operational Data for Run 3 of FY05


written by Sector 7 staff.
Sector 7
Started Jun 9, 2005
(www.mhatt.aps.anl.gov/Sectors/Sector7/Operations/FY05/run3/)



Table of Content:

Introduction.

The work shown below is an account of useful operational data taken during
run 3 of FY05 on 7ID. This page will contain useful stability information on the
7ID High Heat Load Monochromator and also on the various repairs and problems 
identified during the run. It first details the shutdown activities.

Summary of the May 2005 shutdown activities.


Pictures from the May 2005 shutdown.

The cryocooler was serviced this shutdown, and a new roof platform was built to
move the cryocooler in the future to the roof of 7ID-A. This move should occur
in the next shutdown and is motivated by safety, beam stability and space
constraints.

Note that the white beam microdiffraction set up was taken apart in 7ID-B. The
large table inside the hutch was exchanged for a 4-circle diffractometer 
belonging to Sector 8. The 7ID-B in the future will have a 4-circle
diffractometer and an imaging set up for time-resolved phenomena. A new roof
labyrinth was also build to ease the access to the experimental hutch.

7ID-C and 7ID-D experiments now can rely on vacuum flight paths built by Don
Walko! Now all our hutches flight paths can be evacuated to reduce air
absorption and scattering.

One of the Michigan optical table was rewired for use with ACS StepPak movable
motor stand. The limits were fixed as they were reversed and the wiring can now
accept either the NextStep drivers or the ACS StepPak units.

The new laser enclosure was built in May, and will be completed during the run
as time permits. 

A new cable tray was built also to allow several of our controls station to move
to the new partition adjacent to the experimental hall.

There were inumerable numbers of computer upgrades this shutdown, all worth
reading about on  our new computer log book (local access only due to computer 
security).  Note that Dohn Arms has also purchased new LCD screens for all
beamline computers, installed DVD/CD burners on all 3 beamline PCs and much
more.

June 13, 2005.

 Report from Eric Landahl that the mono 
no longer scans properly near the Kr edge.

Useful ELOGs from the summer 2005 run (local access only)

Lessons learned during run 3 of FY05


We should set up separate electronics for the long 200mm KB mirror system.
We keep loosing the parameters when we switch the unit from 7ID-C to 7ID-D.
See ELOG 238 on example of problems associated with switching.

We should also fix the 7ID-D horizontal mirror mount so that the mirror does
not fall down when the horizontal bending forces are released.

U.S. Department of Energy UChicago Argonne LLC Office of Science - Department of EnergyOffice of Basic Energy Sciences - Department of Energy
Privacy & Security Notice | Contact Us | Site Map